The length of one leg of a right triangle is 3 centimeters more than the length of the other leg. The length of the hypotenuse is 15 centimeters. Find the lengths of the two legs.
Answer by oberobic(2304)

Using the Pythagorean Theorem, we know that
c^2 = a^2 + b^2
.
c^2 = 15^2
c^2 = 225
.
a = b+3
.
a^2 + b^2 = 225
(b+3)^2 + b^2 = 225
(b+3)(b+3) + b^2 = 225
Use FOIL on the first term.
b^2 +3b +3b + 9 +b^2 = 225
Collect terms.
2b^2 +6b + 9 = 225
Subtract 225 from both sides
2b^2 +6b -216 = 0
Divide both sides by 2 to eliminate the coefficient on b^2.
b^2 +3b -108 = 0
Factor
(b+12)(b-9) = 0
So, b=-12 or b=9.
A negative length does not make sense, so b=9.
.
Use the Pythagorean to find side a.
.
a^2 = c^2 - b^2
a^2 = 225 - 81
a^2 = 144
a = 12
.
So, the lengths of the two legs are 9 cm and 12 cm.
.
Done.